OpenStudy (anonymous):

whats y^6/y^2 simplified

OpenStudy (bibby):

The general rule for dividing exponents is \[ \frac{ x^a }{ x^b } = x^(a-b)\]

OpenStudy (anonymous):

so would it be y^3?

OpenStudy (bibby):

6-2 is not 3. If you want a different way to visualize things, lemme know if this helps